what is the molarity of a solution containing 408grams of NaOH in 2 Liters of Water? (Round to the 2nd decimal)



1
Expert's answer
2022-03-25T12:21:03-0400

Molar mass of NaOH is:

M(NaOH)=M(Na)+M(O)+M(H)=22.990+15.999+1.008=39.997 (g/mol).

Then 408g is 408/39.997=10.201 (mol).

Molarity of solution is then 10.201/2=5.1005 (mol/L).

ANSWER: 5.1 mol/L